OpenStudy (chefcurry30):

what is the solution to the problem, 2n-7/3=15

sam (.sam.):

Is it \[2n-\frac{7}{3}=15\]?

OpenStudy (chefcurry30):

the 2n is in the fraction to

sam (.sam.):

\[\frac{2n-7}{3}=15\] Any ideas?

sam (.sam.):

You can multiply both sides by 3 to get rid of the denominator. \[\frac{2n-7}{3}=15 \\ \\ 2n-7=45\] Then add 7 and then divide both sides by 2 you should be able to get the answer.

OpenStudy (chefcurry30):

thanks a lot